OpenStudy (anonymous):

Thanks for helping! I need to convert complex numbers into rectangular form. I'm stuck on this one: 5(cos(5pi/4)+isin(5pi/4))

OpenStudy (anonymous):

@Hero @perl @jim_thompson5910 @ganeshie8 @aaronq @zepdrix

OpenStudy (anonymous):

@TheSmartOne @Loser66 @Luigi0210

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

this is not nearly as hard as it looks evaluate the functions and distribute is all

OpenStudy (misty1212):

do you know \(\cos(\frac{5\pi}{4})\)?

OpenStudy (anonymous):

-1/sqrt(2)

OpenStudy (misty1212):

yeah most people write \(-\frac{\sqrt 2}{2}\) but it is the same thing

OpenStudy (anonymous):

I'm sort of confused about the isin though

OpenStudy (misty1212):

how about \(\sin(\frac{5\pi}{4})\)

OpenStudy (anonymous):

Is there a formula I need to follow?

OpenStudy (anonymous):

-sqrt(2)/2

OpenStudy (misty1212):

let me ask you a question how is it you know the sine, but not the cosine?

OpenStudy (anonymous):

?

OpenStudy (misty1212):

or the other way around seems like if you are able to find one, you should be able to find the other it is not like one is more difficult or something

OpenStudy (misty1212):

yeah they are both \(-\frac{\sqrt 2}{2}\)

OpenStudy (anonymous):

so I just plug it in and simplify?

OpenStudy (misty1212):

so \[5\left (-\frac{\sqrt2}{2}-i\frac{\sqrt 2}{2}\right)\]

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

then distribute

OpenStudy (misty1212):

that is all yes

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

thanks!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]